OpenStudy (anonymous):

How do I calculate the absolute uncertainty when there is an exponent? (see question 2)

OpenStudy (anonymous):

1) when the number is (5.3322∗ 10^-5) would the uncertainty be (1*10^-9) or (0.0001)?

OpenStudy (anonymous):

2) how to calculate the absolute uncertainty when there is an exponent? For example \[(-7.5899*10^{-6}) *(22.62 \pm 0.01) \].

OpenStudy (anonymous):

for example lets say u are given 10.1 you dont know if it was 10.05 to 10.15 so your uncertainty is 10.1 +/- 0.05

OpenStudy (anonymous):

We were told in class that when the uncertainty is implicit we will assume its the last digit and it will always be 1. For example: 0.0967 the uncertainty will be 0.0001.

OpenStudy (anonymous):

(5.3322∗ 10^-5) same thing this part 5.3322 you dont know if it was 5.33215 to 5.33225 so 5.3322 +/- 0.00005

OpenStudy (anonymous):

ok i guess that maybe fine too its like roudning up the uncertainty itself

OpenStudy (anonymous):

+/- 0.00005 = +/- 0.0001

OpenStudy (anonymous):

ooohhhh ok. so the uncertainty will be determined regardless of the exponent.

OpenStudy (anonymous):

yeah that is right

OpenStudy (anonymous):

(5.33220∗ 10^-5) if they want to go upto exponent

OpenStudy (anonymous):

you have to add that 0

OpenStudy (anonymous):

ok. thanks!!

OpenStudy (anonymous):

sure np

OpenStudy (anonymous):

for your question 2, absolute uncertainty is that the uncertainty that is based on proportion

OpenStudy (anonymous):

iwhen y=x^n, is the formula n* (relative uncertainty), relative uncertainty= (uncertainty/value)

OpenStudy (anonymous):

okay i see

OpenStudy (anonymous):

here is a more logical approach to calculating your absolute uncertainty

OpenStudy (anonymous):

suppose you have 10.0+/- 0.1 and 2.0+/- 0.2 you see 10.0*2.0 = 20.0 then you see the max and mix that happens with the uncertainty 10.1*2.2 = 20.2 + 2.02 = 22.22 and the min 9.9*1.8 = 18-0.18=17.82

OpenStudy (anonymous):

(22.22-17.82)/2 would be the uncertainty?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me check how u actually find absolute uncertainties hmm

OpenStudy (anonymous):

look at this http://web.uvic.ca/~jalexndr/192UncertRules.pdf

OpenStudy (anonymous):

that is probably best forget what i said so far

OpenStudy (anonymous):

ok. thank you so much!!